# Multiplayer Federated Learning: Reaching Equilibrium with Less Communication

### v1

@@ Abstract
Traditional Federated Learning (FL) approaches assume collaborative clients with aligned objectives working towards a shared global model. However, in many real-world scenarios, clients act as rational players with individual objectives and strategic behaviors, a concept that existing FL frameworks are not equipped to adequately address. To bridge this gap, we introduce Multiplayer Federated Learning (MpFL), a novel framework that models the clients in the FL environment as players in a game-theoretic context, aiming to reach an equilibrium. In this scenario, each player tries to optimize their own utility function, which may not align with the collective goal. Within MpFL, we propose Per-Player Local Stochastic Gradient Descent (PEARL-SGD), an algorithm in which each player/client performs local updates independently and periodically communicates with other players. We theoretically analyze PEARL-SGD and prove that it reaches a neighborhood of equilibrium with less communication in the stochastic setup compared to its non-local counterpart. Finally, we verify our theoretical findings through numerical experiments.
@@ 1 Introduction
Federated Learning (FL) has emerged as a powerful collaborative learning paradigm where multiple clients jointly train a machine learning model without sharing their local data. In the classical FL setting, a central server coordinates multiple clients (e.g., mobile devices, edge devices) to collaboratively learn a shared global model without exchanging their own training data [48, 54, 79, 64]. In this scenario, each client performs local computations on its private data and periodically communicates model updates to the server, which aggregates them to update the global model. This collaborative approach has been successfully applied in various domains, including natural language processing [69, 43], computer vision [70, 63], and healthcare [4, 116].
@@
Despite their success, traditional FL frameworks rely on the key assumption that all participants are fully cooperative and share aligned objectives, collectively working towards optimizing the performance of a shared global model (e.g.,  minimizing the average of individual loss functions). This assumption overlooks situations where participants have individual objectives, or competitive interests that may not align with the collective goal. Diverse examples of such scenarios have been extensively considered in the game theory literature, including Cournot competition in economics [2], optical networks [91], electricity markets [98], energy consumption control in smart grid [120], or mobile robot control [49]. Despite their relevance, these applications have yet to be associated with FL, presenting an unexplored opportunity to bridge game theory and FL for more robust and realistic frameworks.
@@
To address these limitations of classical FL approaches, we propose a novel framework called Multiplayer Federated Learning (MpFL), which models the FL process as a game among rational players with individual utility functions. In MpFL, each participant is considered a player who aims to optimize their own objective while interacting strategically with other players in the network via a central server. This game-theoretic perspective acknowledges that participants may act in their self-interest, have conflicting goals, or be unwilling to fully cooperate. By incorporating these dynamics, MpFL provides a more realistic and flexible foundation for FL in competitive and heterogeneous environments.
@@
In the literature, there are multiple strategies that aim to incorporate a personalization approach into classical FL, including multi-task learning [106, 82], transfer learning [52], and mixing of the local and global models [40, 41], to name a few. However, to the best of our knowledge, none of them is able to formulate the behaviour of the clients/players in a non-cooperative environment. This gap is precisely what Multiplayer Federated Learning (MpFL) aims to address.
@@ 1.1 Main contributions
In this work, we make the following main contributions:
@@
    • Deterministic setting: For the full-batch (deterministic) variant of PEARL-SGD, we prove that under suitable assumptions, PEARL-SGD converges linearly to an equilibrium for any communication period $\tau>1$, provided that the constant step-size $\gamma$ is sufficiently small (see Theorem 3.3).
@@
        Constant step-size: We show that under the same assumptions as in the deterministic case, PEARL-SGD converges linearly to a neighborhood of equilibrium (see Theorem 3.4). In Corollary 3.5, we show that with appropriate step-size depending on the total number of local SGD iterations $T$, PEARL-SGD achieves $\tilde{{\mathcal{O}}}(1/T)$ convergence rate with improved communication complexity when $T$ is sufficiently large.
        Decreasing step-size rule: We prove that PEARL-SGD converges to an exact equilibrium (without neighborhood of convergence) with sublinear convergence (see Theorem 3.6). In this scenario, the asymptotic rate and communication complexity are essentially the same as in Corollary 3.5, but this result does not require the step-sizes to depend on $T$.
  • Numerical Evaluation: We provide extensive numerical experiments verifying our theoretical results and show the benefits in terms of communications of PEARL-SGD over its non-local counterpart in the MpFL settings.
Table 1: Summary of theoretical results for PEARL-SGD. Theorem 3.3 considers the full-batch (deterministic) scenario. Theorem 3.4 and Theorem 3.6 both considers the general stochastic case. These results differ in the step-size choice; the former uses a constant step-size, while the latter uses decreasing step-sizes. In the Convergence column, “Linear” and “Sublinear” indicates the convergence rate, “Exact” refers to convergence to an equilibrium, and “Neighborhood” refers to convergence to a neighborhood of an equilibrium.
@@
Theorem 3.3 Deterministic Constant
Linear$+$Exact
Theorem 3.4 Stochastic
@@
Linear$+$Neighborhood
Theorem 3.6 Stochastic
@@ 2 Multiplayer Federated Learning: Definition and Related Work
In this section, we introduce the Multiplayer Federated Learning (MpFL) framework and explain its main differences compared to the classical FL [48], federated minimax optimization [19, 104, 124] and personalized FL [26, 111].
@@ 2.1 Definition of MpFL
Let $x^{i}\in\mathbb{R}^{d_{i}}$ denote the action of player $i=1,\dots,n$ and let ${\mathbf{x}}=(x^{1},\dots,x^{n})\in\mathbb{R}^{D}=\mathbb{R}^{d_{1}+\cdots+d_{n}}$ be the joint action/strategy vector of all players. Let $f_{i}(x^{1},\dots,x^{n})\colon\mathbb{R}^{d_{1}+\dots+d_{n}}\to\mathbb{R}$ be the function of the player $i$ (which player $i$ prefers to minimize in $x^{i}$) and let $x^{-i}=(x^{1},\dots,x^{i-1},x^{i+1},\dots,x^{n})\in\mathbb{R}^{D-d_{i}}$ be the vector containing all players’ actions except that of player $i$. With this notation in place, the goal in a $n$ player game is to find an equilibrium, a joint action ${\mathbf{x}}_{\star}=(x_{\star}^{1},\dots,x_{\star}^{n})\in\mathbb{R}^{D}$, formally expressed as:
@@
In our proposed FL environment, each client/player uses the strategies of all players to execute local updates. In particular, each player keeps the other players’ strategies fixed and updates their own value, which is later shared with the master server, which concatenates all new strategies and sends them back to all players. Later, in Section 3 we introduce and analyze Algorithm 1, named Per-Player Local SGD (PEARL-SGD), which formalizes the above setting.
Similarly to the classical FL regime, our setting focuses on heterogeneous data (non-i.i.d.) as we do not make any restrictive assumption on the data distribution ${\mathcal{D}}_{i}$ or the similarity between the functions of the players.
@@
Let us present the main assumptions on the functions of the multiplayer game, which we later use to provide the convergence analysis for the proposed Per-Player Local SGD. In our work, we make two main assumptions on the functions $f_{i}$ of each player $i\in[n]$. We assume that the function is convex and smooth.
Throughout this work, we denote the gradient of $f_{i}$ (function of player $i\in[n]$) with respect to $x^{i}$ by:
$\nabla_{x^{i}}f_{i}(x^{1},\dots,x^{n})=\nabla f_{i}(x^{i};x^{-i}).$
This convention allows us to remove the cumbersome subscript $x^{i}$ from the $\nabla$ notation; we only differentiate $f_{i}$ with respect to $x^{i}$ but never with $x^{-i}$.
@@ 2.2 Comparison with closely related FL frameworks
Having presented the MpFL setting, let us now provide a concise survey of related setups: classical FL, federated minimax optimization and personalized FL. We compare each of them with our proposed MpFL. An additional list of related work is provided in Appendix A.
@@
In its basic formulation, classical federated learning can be expressed as the minimization of the objective function [48],
$\begin{array}[]{cc}\underset{x\in\mathbb{R}^{d}}{\textrm{minimize}}&f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\quad\text{where}\quad f_{i}(x)=\mathbb{E}_{\xi^{i}\sim{\mathcal{D}}_{i}}[F_{i}(x,\xi^{i})].\end{array}$
Here, $x\in\mathbb{R}^{d}$ represents the parameter for the global model, $f_{i}$ denotes the local objective function at client $i$, and ${\mathcal{D}}_{i}$ denotes the data distribution of client $i$. The local loss functions $F_{i}(x,\xi^{i})$ are often the same across all clients, but the local data distribution ${\mathcal{D}}_{i}$ will often vary, capturing data heterogeneity. The foundational communication-efficient algorithm for this setup is FedAvg (Local SGD), proposed and massively popularized by McMahan et al. [78]. Despite its simplicity, Local SGD has shown empirical success in terms of convergence speed and communication frequency, and many works have provided theoretical explanation for this performance [109, 23, 110, 51].
In these works, clients work in a fully cooperative manner to find $x_{\star}=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}f(x)$, unlike our proposed MpFL where the clients who now serve as players of the game seek an equilibrium among possibly competing (non-cooperative) objectives.
@@
Federated minimax optimization is a more recent concept proposed as a federated extension of minimax optimization problems appearing in many machine learning applications. In that scenario, the main problem can be expressed as:
$\underset{x\in\mathbb{R}^{d_{x}}}{\textrm{minimize}}\,\,\underset{y\in\mathbb{R}^{d_{y}}}{\textrm{maximize}}\quad{\mathcal{L}}(x)=\frac{1}{n}\sum_{i=1}^{n}{\mathcal{L}}_{i}(x,y)\quad\text{where}\quad{\mathcal{L}}_{i}(x,y)=\mathbb{E}_{\xi^{i}\sim{\mathcal{D}}_{i}}[\phi_{i}(x,y,\xi)].$
Here $n$ is the number of clients, and ${\mathcal{L}}_{i}(x,y)$ represents the local loss function at client $i$ that depends on both $x$ and $y$, defined as ${\mathcal{L}}_{i}(x,y)=\mathbb{E}_{\xi^{i}\sim{\mathcal{D}}_{i}}[\phi_{i}(x,y,\xi)]$. Note that in that scenario, each client has access to the information of both players $x$ and $y$. $\phi_{i}(x,y,\xi)$ denotes the loss for the data point $\xi$, sampled from the local data distribution ${\mathcal{D}}_{i}$ at client $i$. Based on the properties of the model, the functions ${\mathcal{L}}_{i}(x,y)$ can be smooth/non-smooth, convex/non-convex with respect to player $x$, and concave/non-concave with respect to player $y$. The extension of Local SGD for solving this problem are Local Stochastic Gradient Descent-Ascent (SGDA) [19, 104] or Local Stochastic Extragradient (SEG) [6, 7] algorithms. More recently there was also an approach based on primal-dual updates [14].
While this line of work also studied federated learning in the context of minimax optimization and games, it is totally different from MpFL. As we explained above, the setup in this scenario assumes that each FL client has access to both players of the minimax game, and they do not take the multiplayer aspect into account. In contrast, in our setting, we assume that each client is a player of a large-scale multiplayer game who only has access to their objective $f_{i}$ and its gradient and only updates their action $x^{i}$. In our work, we design the novel PEARL-SGD algorithm, which is suitable for the MpFL setting, a task not possible using the existing Local SGDA and Local SEG methods.
@@
In personalized FL [26, 111, 41, 40, 20, 112], clients aim to learn models tailored to each local data distribution, while generalizing well over the data from other clients [42]. One way of formulating personalized FL is [41, 42]:
$\underset{{\mathbf{x}}=(x^{1},\dots,x^{n})\in\mathbb{R}^{nd}}{\textrm{minimize}}\quad\frac{1}{n}\sum_{i=1}^{n}h_{i}(x^{i})+\frac{\lambda}{2n}\sum_{i=1}^{n}\left\|x^{i}-\overline{x}\right\|^{2}$
where $x^{1},\dots,x^{n}\in\mathbb{R}^{d}$ are local models of each player, $h_{1},\dots,h_{n}$ are the local losses following the each player’s distributions, $\overline{x}=\frac{1}{n}\sum_{i=1}^{n}x^{i}$ is the average model and $\lambda>0$ is regularization strength. Given that each $h_{i}$ is convex, the first-order optimality condition for the above is $x^{i}_{\star}-\overline{x}_{\star}+\frac{1}{\lambda}\nabla h_{i}(x^{i}_{\star})=0$ [41]. This is equivalent to the equilibrium condition for the $n$-player game (MpFL) where each player has the objective function $f_{i}(x^{i};x^{-i})=h_{i}(x^{i})+\frac{\lambda}{2}\sum_{i=1}^{n}\left\|x^{i}-\overline{x}\right\|^{2}$. In this respect, personalized FL can be viewed as a particular instance of MpFL.

### v2

@@ Abstract
Traditional Federated Learning (FL) approaches assume collaborative clients with aligned objectives working toward a shared global model. However, in many real-world scenarios, clients act as rational players with individual objectives and strategic behavior, a concept that existing FL frameworks are not equipped to adequately address. To bridge this gap, we introduce Multiplayer Federated Learning (MpFL), a novel framework that models the clients in the FL environment as players in a game-theoretic context, aiming to reach an equilibrium. In this scenario, each player tries to optimize their own utility function, which may not align with the collective goal. Within MpFL, we propose Per-Player Local Stochastic Gradient Descent (PEARL-SGD), an algorithm in which each player/client performs local updates independently and periodically communicates with other players. We theoretically analyze PEARL-SGD and prove that it reaches a neighborhood of equilibrium with less communication in the stochastic setting compared to its non-local counterpart. Finally, we verify our theory through numerical experiments.
@@ 1 Introduction
Federated Learning (FL) has emerged as a powerful collaborative learning paradigm where multiple clients jointly train a machine learning model without sharing their local data. In the classical FL setting, a central server coordinates multiple clients (e.g., mobile or edge devices) to collaboratively learn a shared global model without exchanging their own training data [kairouz2021advances, konevcny2016federated, mcmahan2017communication, li2020federated]. In this scenario, each client performs local computations on its private data and periodically communicates model updates to the server, which aggregates them to update the global model. This collaborative approach has been successfully applied in various domains, including natural language processing [liu2021federated, hard2018federated], computer vision [LiuHuangLuoHuangLiuChenFengChenYuYang2020_fedvision, li2021survey], and healthcare [antunes2022federated, xu2021federated].
@@
Despite their success, traditional FL frameworks rely on the key assumption that all participants are fully cooperative and share aligned objectives, collectively working to optimize the performance of a shared global model (e.g., minimizing the average of individual loss functions). This assumption overlooks situations where participants have individual objectives or competitive interests that may not align with the collective goal. A variety of such scenarios have been extensively considered in the game theory literature, including Cournot competition in economics [AhmedAgiza1998_dynamics], optical networks [PanPavel2007_global], electricity markets [SaadHanPoorBasar2012_gametheoretic], energy consumption control in smart grid [YeHu2017_game], or mobile robot control [KalyvaPsillakis2024_distributed]. These applications have yet to be associated with FL, presenting an underexplored opportunity to bridge game theory and FL for more robust and realistic frameworks.
@@
To address these limitations of classical FL approaches, we propose a novel framework, Multiplayer Federated Learning (MpFL), which models the FL process as a game among rational players with individual utility functions. In MpFL, each participant is considered a player who aims to optimize their own objective while interacting strategically with other players in the network via a central server. This game-theoretic perspective acknowledges that participants may act in their self-interest, have conflicting goals, or be unwilling to fully cooperate. By incorporating these dynamics, MpFL provides a more realistic and flexible foundation for FL in competitive and heterogeneous environments.
@@
In the literature, there are multiple strategies that aim to incorporate personalization into classical FL, including multi-task learning [SmithChiangSanjabiTalwalkar2017_federated, mills2021multi], transfer learning [KhodakBalcanTalwalkar2019_adaptive], and mixing of the local and global models [hanzely2020federated, HanzelyHanzelyHorvathRichtarik2020_lower], to name a few. However, to the best of our knowledge, none of them can formulate the behavior of the clients/players in a non-cooperative environment. This gap is precisely what Multiplayer Federated Learning (MpFL) aims to address.
@@ 1.1 Main contributions
    • Deterministic setting: For the full-batch (deterministic) variant of PEARL-SGD, we prove that under suitable assumptions, PEARL-SGD converges linearly to an equilibrium for any communication period $\tau>1$, provided that the constant step-size $\gamma$ is sufficiently small (see Theorem˜3.3).
@@
        Constant step-size: We show that under the same assumptions as in the deterministic case, PEARL-SGD converges linearly to a neighborhood of equilibrium (see Theorem˜3.4). In Corollary˜3.5, we show that with appropriate step-size depending on the total number of local SGD iterations $T$, PEARL-SGD achieves $\tilde{{\mathcal{O}}}(1/T)$ convergence rate with improved communication complexity when $T$ is sufficiently large.
        Decreasing step-size rule: We prove that PEARL-SGD converges to an exact equilibrium (without neighborhood of convergence) with sublinear convergence (see Theorem˜3.6). In this scenario, the asymptotic rate and communication complexity are essentially the same as in Corollary˜3.5, but this result does not require the step-sizes to depend on $T$.
  • Numerical Evaluation: We provide numerical experiments verifying our theoretical results and show the benefits in terms of communications of PEARL-SGD over its non-local counterpart in the MpFL settings.
Table 1: Summary of theoretical results for PEARL-SGD. Theorem˜3.3 considers the full-batch (deterministic) scenario. Theorem˜3.4 and Theorem˜3.6 both considers the general stochastic case. These results differ in the step-size choice; the former uses a constant step-size, while the latter uses decreasing step-sizes. In the Convergence column, “Linear” and “Sublinear” indicates the convergence rate, “Exact” refers to convergence to an equilibrium, and “Neighborhood” refers to convergence to a neighborhood of an equilibrium.
@@
Theorem˜3.3 Deterministic Constant
Linear++Exact
Theorem˜3.4 Stochastic
@@
Linear++Neighborhood
Theorem˜3.6 Stochastic
@@ 2 Multiplayer Federated Learning: Definition and Related Work
In this section, we introduce the Multiplayer Federated Learning (MpFL) framework and explain its main differences compared to the classical FL [kairouz2021advances], federated minimax optimization [deng2021local, sharma2022federated, zhang2023communication] and personalized FL [FallahMokhtariOzdaglar2020_personalized, T.DinhTranNguyen2020_personalized].
@@ 2.1 Definition of MpFL
Let $x^{i}\in\mathbb{R}^{d_{i}}$ denote the action of player $i=1,\dots,n$ and let ${\mathbf{x}}=(x^{1},\dots,x^{n})\in\mathbb{R}^{D}=\mathbb{R}^{d_{1}+\cdots+d_{n}}$ be the joint action/strategy vector of all players. Let $f_{i}(x^{1},\dots,x^{n})\colon\mathbb{R}^{d_{1}+\dots+d_{n}}\to\mathbb{R}$ be the function of the player $i$ (which player $i$ prefers to minimize in $x^{i}$) and let $x^{-i}=(x^{1},\dots,x^{i-1},x^{i+1},\dots,x^{n})\in\mathbb{R}^{D-d_{i}}$ be the vector containing all players’ actions except that of player $i$. With this notation in place, the goal of an $n$-player game is to find an equilibrium, a joint action ${\mathbf{x}}_{\star}=(x_{\star}^{1},\dots,x_{\star}^{n})\in\mathbb{R}^{D}$, formally expressed as:
@@
In our proposed FL environment, each client/player uses the strategies of all players to execute local updates. In particular, each player keeps the other players’ strategies fixed and updates their own value, which is later shared with the master server, which concatenates all new strategies and sends them back to all players. Later, in Section 3 we introduce and analyze Algorithm˜1, named Per-Player Local SGD (PEARL-SGD), which formalizes the above setting.
Similarly to classical FL, our setting allows for heterogeneous (non-iid) data as we have no restrictive assumption on the data distribution ${\mathcal{D}}_{i}$ or the similarity between the functions of the players.
@@
Let us present the main assumptions on the functions of the multiplayer game, which we later use to provide the convergence analysis for PEARL-SGD—objective function $f_{i}$ of each player $i\in[n]$ is convex and smooth.
Throughout this work, we denote the partial derivative of $f_{i}$ with respect to $x^{i}$ as $\nabla_{x^{i}}f_{i}(x^{1},\dots,x^{n})=\nabla f_{i}(x^{i};x^{-i})$. This convention allows us to remove the cumbersome subscript $x^{i}$ from the $\nabla$ notation, with the understanding that we only differentiate $f_{i}$ with respect to $x^{i}$ but never with $x^{-i}$.
@@ 2.2 Comparison with closely related FL frameworks
Having presented the MpFL setting, let us now provide a concise survey of the most closely related setups: classical FL, federated minimax optimization and personalized FL. We compare each of them with MpFL, and explain in detail in Appendix˜B why existing FL algorithms are not applicable to MpFL. An additional list of related work, explaining the literatures on game theory, distributed Nash equilibrium search, learning in games and the usage of game theory for modeling clients’ social behavior in FL, and how MpFL is distinguished from them, is provided in Appendix˜A.
@@
The basic formulation for classical FL is [kairouz2021advances]: $\underset{x\in\mathbb{R}^{d}}{\textrm{minimize}}\,\,f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)$. Here, $x\in\mathbb{R}^{d}$ represents the global model parameter, $f_{i}(x)=\mathbb{E}_{\xi^{i}\sim{\mathcal{D}}_{i}}[F_{i}(x,\xi^{i})]$ denotes the local objective function at client $i$, and ${\mathcal{D}}_{i}$ denotes the data distribution of client $i$. The local loss functions $F_{i}(x,\xi^{i})$ are often of the same form across all clients, but the data distribution ${\mathcal{D}}_{i}$ generally varies, capturing data heterogeneity. The foundational communication-efficient algorithm for this setup is FedAvg (Local SGD), proposed and popularized by [McMahanMooreRamageHampsonArcas2017_communicationefficient]. Despite its simplicity, Local SGD has shown empirical success in terms of convergence speed and communication cost. Many works provided theoretical explanation for this performance [Stich2019_local, DieuleveutPatel2019_communication, StichKarimireddy2020_errorfeedback, KhaledMishchenkoRichtarik2020_tighter].
Note that in these works on classical FL, clients work in a fully cooperative manner to find $x_{\star}=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}f(x)$, unlike in our proposed MpFL, where the clients serve as players of the game and seek an equilibrium among possibly competing (non-cooperative) objectives.
@@
This is a more recent, federated extension of minimax optimization appearing in many ML applications. There the problem is: $\underset{x\in\mathbb{R}^{d_{x}}}{\textrm{minimize}}\,\,\underset{y\in\mathbb{R}^{d_{y}}}{\textrm{maximize}}\,\,{\mathcal{L}}(x,y)=\frac{1}{n}\sum_{i=1}^{n}{\mathcal{L}}_{i}(x,y)$. Here $n$ is the number of clients and ${\mathcal{L}}_{i}$ represents the local loss function at client $i$ that depends on both $x$ and $y$. It is defined as ${\mathcal{L}}_{i}(x,y)=\mathbb{E}_{\xi^{i}\sim{\mathcal{D}}_{i}}[\phi_{i}(x,y,\xi)]$, where $\phi_{i}(x,y,\xi)$ denotes the loss for the data point $\xi$, sampled from the local data distribution ${\mathcal{D}}_{i}$ at client $i$. The extension of Local SGD for solving this problem are Local Stochastic Gradient Descent-Ascent (SGDA) [deng2021local, sharma2022federated] or Local Stochastic Extragradient (SEG) [beznosikov2020distributed, beznosikov2022decentralized] algorithms. More recently there was also an approach based on primal-dual updates [CondatRichtarik2022_randprox].
While this line of work also studied federated learning in the context of minimax optimization and games, it is totally different from MpFL. The setup assumes that each FL client adjusts the actions of both players $x$ and $y$ of the minimax game, and does not take the multiplayer aspect into account. In contrast, in MpFL, we assume that each client $i$ is a player of a large-scale multiplayer game who only adjusts their own action $x^{i}$ in the interest of optimizing their own objective $f_{i}$. In our work, we design the novel PEARL-SGD algorithm suitable for MpFL, as the existing Local SGDA and Local SEG methods cannot handle the setup.
@@
In personalized FL [FallahMokhtariOzdaglar2020_personalized, T.DinhTranNguyen2020_personalized, HanzelyHanzelyHorvathRichtarik2020_lower, hanzely2020federated, DengKamaniMahdavi2020_adaptive, TanYuCuiYang2023_personalized], clients aim to learn models tailored to each local data distribution, while generalizing well on all clients’ data [HanzelyZhaokolar2023_personalized]. One formulation of personalized FL is $\underset{{\mathbf{x}}=(x^{1},\dots,x^{n})\in\mathbb{R}^{nd}}{\textrm{minimize}}\,\,\frac{1}{n}\sum_{i=1}^{n}h_{i}(x^{i})+\phi(x^{1},\dots,x^{n})$; in [HanzelyHanzelyHorvathRichtarik2020_lower, HanzelyZhaokolar2023_personalized], for example, $\phi$ is taken as the model consensus regularizer $\phi(x^{1},\dots,x^{n})=\frac{\lambda}{2n}\sum_{i=1}^{n}\left\|x^{i}-\overline{x}\right\|^{2}$. Given that each $h_{i}$ and $\phi$ are convex, its first-order optimality condition is equivalent to the equilibrium condition for the $n$-player game where players’ objectives are $f_{i}(x^{i};x^{-i})=h_{i}(x^{i})+n\phi(x^{1},\dots,x^{n})$. Such formulation of personalized FL, therefore, is an instance of MpFL. On the other hand, there exist personalized FL approaches that are not modeled as a direct subclass of MpFL—e.g., [FallahMokhtariOzdaglar2020_personalized, T.DinhTranNguyen2020_personalized] proposed to maintain a separate global model $w$ along with local models $x^{i}$. Still, it is clear that MpFL is closely related to personalized FL. Importantly, however, the purpose of MpFL is not limited to having personalized models suitable for local data distribution, and encompasses settings where $x^{1},\dots,x^{n}$ differ in dimensionality or structure.